# FedCCRL: Federated Domain Generalization with Cross-Client Representation Learning

### v1

@@ 3.4 Representation Alignment
$\mathcal{L}_{\text{SC}}(Z^{\prime},Z^{\prime\prime})=\sum_{i\in I}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{s(Z_{i},Z_{p})}{\sum_{a\in A(i)}s(Z_{i},Z_{a})},$ (11)
@@
$\mathcal{L}_{\text{RA}}=\frac{1}{2}(\mathcal{L}_{SC}(Z^{(1)},Z)+\mathcal{L}_{SC}(Z^{(2)},Z)).$ (12)
3.5 Jensen-Shannon Divergence
As $X^{(1)}$ and $X^{(2)}$ are generated from $X$, we expect the model to not only correctly classify the original and augmented samples, but also produce similar predictions $\hat{Y}$, $\hat{Y}^{(1)}$ and $\hat{Y}^{(2)}$. We utilize Jensen-Shannon divergence [20] minimization to achieve consistent predictions across original and generated samples. The Jensen-Shannon divergence is designed to measure the similarity between these probability distributions by averaging them and then computing their divergence from the mean. Specifically, let the average prediction distribution be defined as:
$\bar{Y}=\frac{1}{3}(\hat{Y}+\hat{Y}^{(1)}+\hat{Y}^{(2)}).$ (13)
The Jensen-Shannon divergence is then computed as follows:
@@ 3.5 Jensen-Shannon Divergence
where $\text{KL}(\cdot,\cdot)$ denotes KL divergence.
@@ 3.6 Overall Loss Function
and $\mathcal{L}_{CRE}$ denotes the cross-entropy loss. The terms $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that control the relative importance of the representation alignment loss $\mathcal{L}_{RA}$ and the Jensen-Shannon divergence loss $\mathcal{L}_{JS}$, respectively.
@@ 4.2 Baselines and Implementation Details
Following the data partition scheme outlined in Section 3.1, we evaluate FedCCRL and the baselines across different numbers of clients. We utilize MobileNetV3-Large [9] as the backbone network, where the final fully connected layer acts as the classifier $g$, and the preceding layers serve as the representation encoder $h$.
For training, we set the number of communication rounds to 10, and the number of local training epochs is fixed at 3 for all methods. The Adam optimizer [13] is employed with an initial learning rate of 0.001. Furthermore, MixStyle is configured with the parameter $\alpha=0.1$, while for AugMix, the parameter $\beta$ is set to 1.0. For supervised contrastive loss, we set the temperature parameter $\tau$ to 0.1. Input samples from the PACS and OfficeHome datasets are resized to $224\times 224$, whereas those from miniDomainNet are resized to $128\times 128$. To manage the learning rate, we adopt a cosine learning rate scheduler, which facilitates a smooth decay of the learning rate throughout the training process.
@@ 4.4 Ablation Study
MixStyle AugMix $\mathcal{L}_{RA}$ $\mathcal{L}_{JS}$ PACS OfficeHome miniDomainNet
@@
We investigate the effects of MixStyle, AugMix, representation alignment regularizer $\mathcal{L}_{RA}$, and Jensen-Shannon divergence $\mathcal{L}_{JS}$ on FedCCRL across the PACS, OfficeHome, and miniDomainNet datasets. The results are presented in Table 2, from which we can draw the following conclusions:
  • All four components significantly enhance model generalization. Using representation alignment loss or Jensen-Shannon divergence individually outperforms cross-entropy alone, and their combination yields the best performance. Additionally, combining AugMix and MixStyle improves results compared to using either technique alone.
  • Even without regularization, FedCCRL with only data augmentation techniques outperforms all baselines. As shown in Table 1, the combination of MixStyle and AugMix surpasses all baselines across the three datasets.
@@ 5 Conclusion
In this paper we propose FedCCRL, a federated domain generalization framework that improves model generalization across unseen domains without violating privacy rules or introducing excessive computational and communication costs. Specifically, FedCCRL effectively combines MixStyle and AugMix for cross-client style transfer and the perturbation of domain-invariant features. In addition, FedCCRL utilizes supervised contrastive loss for representation alignment and apply Jensen-Shannon divergence for prediction alignment. Experiments on PACS, OfficeHome, and miniDomainNet demonstrate that FedCCRL outperforms existing baselines, demonstrating its effectiveness in enhancing models’ generalization capability.

### v3

@@ 3.4 Representation Alignment
$\mathcal{L}_{SC}(Z^{\prime},Z^{\prime\prime})=\sum_{i\in I}\frac{-1}{|P(i)|}\sum_{p\in P(i)}\log\frac{s(Z_{i},Z_{p})}{\sum_{a\in A(i)}s(Z_{i},Z_{a})},$ (11)
@@
$\mathcal{L}_{RA}=\frac{1}{2}(\mathcal{L}_{SC}(Z^{(1)},Z)+\mathcal{L}_{SC}(Z^{(2)},Z)).$ (12)
3.5 Prediction Alignment
Given that $X^{(1)}$ and $X^{(2)}$ are generated from $X$, we seek for the model to accurately classify both the original and augmented samples while also ensuring consistency in the predictions, represented by $\hat{Y}$, $\hat{Y}^{(1)}$, and $\hat{Y}^{(2)}$.
To promote consistent predictions across the original and generated samples, we employ Jensen-Shannon (JS) divergence minimization. The JS divergence is specifically constructed to measure similarity between probability distributions by calculating their average and then measuring each distribution’s divergence from this mean distribution. Formally, let the average prediction distribution be defined as follows:
$\bar{Y}=\frac{1}{3}(\hat{Y}+\hat{Y}^{(1)}+\hat{Y}^{(2)}),$ (13)
where $\bar{Y}$ represents the mean distribution of the predictions. The JS divergence loss is then computed as:
@@ 3.5 Prediction Alignment
where $\text{KL}(\cdot,\cdot)$ denotes KL divergence. Minimizing $\mathcal{L}_{\text{JS}}$ aligns the model’s predictions across original and augmented samples, thus enhancing prediction consistency.
@@ 3.6 Overall Loss Function
and $\mathcal{L}_{CRE}$ denotes the cross-entropy loss. The terms $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters that control the relative importance of the representation alignment loss $\mathcal{L}_{RA}$ and the JS divergence loss $\mathcal{L}_{JS}$ respectively.
3.7 Analysis of Privacy and Additional Costs
Notably, the computational and transmission overhead associated with sample statistics is negligible compared to the costs of model training and transmission. Additionally, attackers cannot reconstruct samples from these statistics or infer model data distribution on each client, thus enhancing privacy protection.
Furthermore, as shown in Figure 6, our method achieves better performance than other baselines even with a minimal upload ratio $r$, further reducing the computational and transmission overhead as well as the risk of privacy leakage.
@@ 4.2 Baselines and Implementation Details
Following the data partition scheme outlined in Section 3.1, we evaluate FedCCRL and the baselines across scenarios with different numbers of clients. We utilize MobileNetV3-Large [9] as the backbone network, where the final fully connected layer acts as the classifier $g$, and the preceding layers serve as the representation encoder $h$.
For training, we set the number of communication rounds to 10, and the number of local training epochs is fixed at 3 for all methods. The Adam optimizer [12] is employed with an initial learning rate of 0.001. Furthermore, CCDT is configured with the parameter $\alpha=0.1$, while for DIFP, the parameter $\beta$ is set to 1.0. For supervised contrastive loss, we set the temperature parameter $\tau$ to 0.1. Input samples from the PACS and OfficeHome datasets are resized to $224\times 224$, whereas those from miniDomainNet are resized to $128\times 128$. To manage the learning rate, we adopt a cosine learning rate scheduler, which facilitates a smooth decay of the learning rate throughout the training process.
@@ 4.4 Ablation Study
CCDT DIFP $\mathcal{L}_{RA}$ $\mathcal{L}_{JS}$ PACS OfficeHome miniDomainNet
@@
We investigate the effects of CCDT, DIFP, representation alignment regularizer $\mathcal{L}_{RA}$, and JS divergence loss $\mathcal{L}_{JS}$ on FedCCRL across the PACS, OfficeHome, and miniDomainNet datasets. The results are presented in Table 2, from which we can draw the following conclusions:
  • All four components significantly enhance the model’s generalization ability. The utilization of representation alignment loss or JS divergence loss independently demonstrates superior performance compared to employing cross-entropy loss in isolation, while their combined application yields the most optimal results. Additionally, combining CCDT and DIFP improves performance compared to using either method alone.
  • Even without presentation alignment or prediction alignment, FedCCRL achieved accuracies of 80.48%, 67.28%, and 60.4% on the PACS, OfficeHome, and miniDomainNet datasets, respectively, outperforming all other baselines.
@@ 5 Conclusion
This paper introduces FedCCRL, a FDG framework that significantly enhances model’s generalization ability while upholding privacy and incurring minimal computational and communication costs. By employing a lightweight cross-client feature extension module, FedCCRL effectively mitigates the challenges posed by limited local data quantity and domain diversity. Furthermore, FedCCRL enables model to capture domain-invariant features with representation and prediction dual-stage alignment.